OpenStudy (anonymous):

Suppose \(\sum_{n=1}^{\infty}a_{n}\) converges, where \(a_{n} \ge 0\ \forall\ n \in \mathbb{N} \ \) Prove that \(\sum_{n=1}^{\infty}a_{n}^{p}\) converges for all \(p > 1\)

OpenStudy (anonymous):

I'm guessing I have to construct some sort of comparison, which means I would have to show \(a_{n} \ge\ a_{n}^{p}\) in order to show \(a_{n}^{p}\) converges. Just not sure how to do that or if my idea is even correct.

OpenStudy (zarkon):

if the sum converges then there exists an \(N\) such that \(a_n<1\) for all \(n\ge N\) thus for \(n\ge N\) we have \(a_n^p<a_n\) when \(p>1\)

ganeshie8 (ganeshie8):

nice :)

OpenStudy (anonymous):

That claim might definitely be enough. I think I can use that and construct something, thanks ^_^

OpenStudy (zarkon):

use the fact that a necessary condition for \(\sum_{n=1}^{\infty}a_{n}\) to converge is that \(a_n\to 0\) as \(n\to\infty\)

OpenStudy (anonymous):

That's exactly where I was going with it, too! Thanks for the confirmation of that :)